OpenStudy (anonymous):

I just need the factors. 8p^2 + 40pq + 50pq^2

OpenStudy (anonymous):

2 can go into these numbers for sure...

OpenStudy (anonymous):

2p(4p+25q^2+20q)

OpenStudy (anonymous):

yeah - always take out the GCF before factoring , as anhhuyalex did above. 8p^2 + 40pq + 50pq^2 =2p( 4p + 20q + 25q^2 ) =2p( 25q^2 + 20q + 4p ) Factor this (and dont forget that what you get needs to be multiplied by that 2p): 25q^2 + 20q + 4p

OpenStudy (anonymous):

Factors = 2, 50

OpenStudy (anonymous):

the factors should add up to 20

OpenStudy (anonymous):

2, 10

OpenStudy (anonymous):

multiplying the numbers should give 100 adding them should give 20

OpenStudy (anonymous):

10, 10

OpenStudy (anonymous):

yes... but something is messed up a bit because of that "p" we have here ..

OpenStudy (anonymous):

OK - Are you sure this is the original question: 8p^2 + 40pq + 50pq^2 Are you sure this is the original question - coz I suspect it might be this: 8p^2 + 40pq + 50q^2 ?

OpenStudy (anonymous):

bit of copy-paste snafoo above - but it should be clear :)

OpenStudy (anonymous):

The original equation is as follows. 8p^2 + 40pq + 50q^2

OpenStudy (anonymous):

OK makes sense now :) That's not what was posted earlier ... phew ...

OpenStudy (anonymous):

oh-no... I wrote the 1 at the beginning of the question wrong.

OpenStudy (anonymous):

OK - so we need to factor the trinomial inside the brackets below: 2( 4p^2 + 20pq + 25q^2 )

OpenStudy (anonymous):

10, 10

OpenStudy (anonymous):

notice, we have: a^2+2ab+b^2=(a+b)^2

OpenStudy (anonymous):

I see it now - It's a Perfect Square. So i need to square root it...